# Ramification filtration via deformations, II

### v3

@@ 1.4. Explicit description of ideals $\mathcal{L}(w)_{W_{M}(k)}$
Set $M(\bar{a},s)=\min\{m\in\mathbb{Z}_{\geqslant 0}\ |\ p^{m}[D_{\bar{a}\bar{n}}]\in\mathcal{L}(r(\bar{a})+s)_{W_{M}(k)}\}$.
We always assume that $r=r(\bar{a})<p$ because $D_{\bar{a}\bar{n}}\in\mathcal{L}(r)_{W_{M}(k)}$.
For $1\leqslant i\leqslant r$, consider the following decompositions in the “base $p$” with $p$-digits from $[0,p)$:
[…]
Set $\bar{u}=u(\bar{a})=\min\{\ j\ |\ (\theta_{j}(a_{1}),\ldots,\theta_{j}(a_{r}))\neq\bar{0}\}$.
Set $\bar{\kappa}_{0}=\kappa_{0}(\bar{a})=\sum_{1\leqslant i\leqslant r}\theta_{\bar{u}}(a_{i})$ and $\bar{\kappa}_{1}=\kappa_{1}(\bar{a})=\sum_{1\leqslant i\leqslant r}\theta_{\bar{u}+1}(a_{i})$.
@@
If $\bar{u}=0$ then $[D_{\bar{a}\bar{n}}]\in\mathcal{L}(r+\bar{\kappa}_{0})_{W_{M}(k)}$.
@@ 1.5. $\mathbb{Z}_{p}$-submodules $\mathcal{L}_{\mathcal{N}}(w)\subset\mathcal{L}(w)$
For $w\geqslant 1$, let $\mathcal{L}_{\mathcal{N}}(w)$ be the closed submodule in $\mathcal{L}$ such that $\mathcal{L}_{\mathcal{N}}(w)_{W_{M}(k)}$ is generated by all $p^{m}[D_{\bar{a}\bar{n}}]$ which satisfy the inequality $|\bar{a}|\geqslant v_{0}U(w-1,m)$.
@@
If $p^{m}[D_{\bar{a}_{1}\bar{n}_{1}},D_{\bar{a}_{2}\bar{n}_{2}}]\in\mathcal{L}_{\mathcal{N}}(w)_{W_{M}(k)}$ then for $i=1,2$, there are $m_{i},w_{i}$ such that:
@@
We can assume that (otherwise, $w_{1}+w_{2}$ is not maximal) for $i=1,2$,
$v_{0}U(w_{i},m_{i})>|\bar{a}_{i}|\geqslant v_{0}U(w_{i}-1,m_{i})\,.$
Let $v_{0}U(w_{1}+w_{2},m)>|\bar{a}_{1}|+|\bar{a}_{2}|$. Then $U(w_{1}+w_{2},m)>U(w-1,m)$ implies $w_{1}+w_{2}\geqslant w$, q.e.d.
@@
Let $m_{1}^{\prime}\geqslant m_{1}$ and $m_{2}^{\prime}\leqslant m_{2}$. Then $v_{0}U(w_{1},m_{1}^{\prime})\geqslant v_{0}U(w_{1},m_{1})>|\bar{a}_{1}|$ and $v_{0}U(w_{2},m_{2}^{\prime})\leqslant v_{0}U(w_{2},m_{2})$.
@@
Clearly, ¯$\bar{\mathcal{L}}(p)=0$ and ¯$\bar{\mathcal{L}}_{W_{M}(k)}$ is generated by $D_{0}$ and all $D_{an}$ with $a<(p-1)v_{0}$.
@@ 2.1. Parameters $r^{*}$ and $N^{*}$
For $1\leqslant u\leqslant u^{*}$, introduce the following subset in $\mathbb{Q}$ :
$A[u]$ is the set of all
$a_{1}p^{n_{1}}+a_{2}p^{n_{2}}+\dots+a_{u}p^{n_{u}}\,,$
where $n_{1},\dots,n_{u}\leqslant 0$, $n_{1}=0$ and all $a_{i}\in[0,w^{*})\cap\mathbb{Z}$. For $m\in\mathbb{Z}_{\geqslant 0}$, denote by $A[u,m]$ the subset of $A[u]$ consisted of the elements such that all $n_{u}\geqslant-m$.
For $1\leqslant s<p$, $0\leqslant m<M$, let $\delta_{0}(s,m)$ be the minimum of positive values of $v_{0}-\alpha/U(s,m)$ where $\alpha\in A[u,m]$ and $u\leqslant u^{*}$. The existence of such $\delta_{0}(s,m)$ can be proved easily by induction on $u$ for any fixed $s$ and $m$.
@@
For $1\leqslant u\leqslant u^{*}$, introduce the following subset in $\mathbb{Q}$ :
$B[u]$ is the set of all numbers
$r^{*}(b_{1}p^{m_{1}}+b_{2}p^{m_{2}}+\dots+b_{u}p^{m_{u}})\,,$
where all $m_{1},\ldots,m_{u}\leqslant M-1$, $m_{1}\geqslant 0$, $b_{i}\geqslant 0$, $b_{1}\neq 0$, $b_{1}+\dots+b_{u}<p$. (In particular, $0\notin B[u]$.) For $m\in\mathbb{Z}_{\geqslant 0}$, $B[u,m]$ is the subset of $B[u]$ consisted of the elements such that all $m_{u}\geqslant-m$.
The above sets $A[u]$ and an analog of the set $B_{1}[u]:=p^{-(M-1)}B[u]$ (where we required $m_{1}=0$ instead of $-M<m_{1}\leqslant 0$) were introduced in [12]. Proceeding similarly to [12] we obtain that for any $u\leqslant u^{*}$, $A[u]\cap B_{1}[u]=\emptyset$ and there is $\rho(u^{*})>0$ such that
$\rho(A[u],B_{1}[u]):=\inf\{|\alpha-\beta|\ |\ \alpha\in A[u],\beta\in B_{1}[u]\}\geqslant\rho(u^{*})\,.$
@@
a) Use condition C2). Let $\beta_{1}=\beta/p^{M-1}\in B_{1}[u^{*}]$. Then
$q\,|q\alpha-(q-1)\beta_{1}|=q^{2}|\alpha-\beta_{1}+\beta_{1}/q|\geqslant q^{2}|\alpha-\beta_{1}|-\beta q\geqslant q^{2}\rho(A[u^{*}],\,B_{1}[u^{*}])$
$-r^{*}(p-1)q\geqslant 2r^{*}(p-1)q-r^{*}(p-1)q=r^{*}(p-1)q>b^{*}(p-1)\,.$
It remains to multiply this inequality by $p^{M-1}$;
b) Use condition C3): $\alpha/U(s,m)<v_{0}$ implies that $\alpha/U(s,m)<v_{0}-\delta_{0}$ and $q\alpha/U(s,m)<q(v_{0}-\delta_{0})<qr^{*}(1-q^{-1})=b^{*}$.
@@ 2.2. The set $\mathfrak{A}^{0}$
b) Any $p^{m}\alpha$ appears in the form $a_{1}p^{n_{1}}+\ldots+a_{u^{*}}p^{n_{u^{*}}}$, where all $a_{i}\in[0,(p-1)v_{0})$ and $0\leqslant n_{i}<N^{*}$.
c) All $p^{m}\beta$ appear in the form $b^{*}(b_{1}p^{m_{1}}+\ldots+b_{u^{*}}p^{m_{u^{*}}})$, where all $m_{i}\geqslant 0$ and $11\leqslant b_{1}+\ldots+b_{u^{*}}\leqslant p-$ 1.
@@
Suppose $\iota=p^{m}(q\alpha-(q-1)\beta)\in\mathfrak{A}^{0}$. Then:
@@
c) the integers $p^{m}\alpha$ and $p^{m}\beta/r^{*}$ do not depend on the presentation of $\iota$ in the form $p^{m}(qp^{M-1}\alpha-(q-1)\beta)$ from the definition of $\mathfrak{A}^{0}$.
@@
a) If $\iota\in\mathfrak{A}^{0}_{0}$ then $\iota=qp^{M-1}(p^{m}\alpha)\in\mathfrak{A}^{0}_{0}\subset\mathfrak{A}^{0}$ means that
$p^{m}\alpha/(p-1)\leqslant b^{*}/q=r^{*}(1-q^{-1})\in(v_{0}-\delta_{0},v_{0})\,.$
Note that $U(p-1,0)=p-1$ and $\delta_{0}\leqslant\delta(p-1,0)$, cf. Sect.2.1. Threfore, the inequalities $p^{m}\alpha/(p-1)<v_{0}$ and $p^{m}\alpha/(p-1)\leqslant v_{0}-\delta_{0}$ are equivalent. As a result,
$\mathfrak{A}^{0}_{0}\subset\{qp^{M-1}a\ |\ a\in[0,(p-1)v_{0})\cap\mathbb{Z}\}\,.$
The opposite embedding is obvious.
@@
c) If $\iota=p^{m^{\prime}}(qp^{M-1}\alpha^{\prime}-(q-1)\beta^{\prime})$ is another presentation of $\iota$ then $p^{m}\beta/r^{*}$ and $p^{m^{\prime}}\beta^{\prime}/r^{*}$ are non-negative congruent modulo $qp^{M-1}$ integers and the both are smaller than $qp^{M-1}$. Indeed, we have
$\beta/r^{*}=b_{1}p^{m_{1}}+b_{2}p^{m_{2}}+\dots+b_{u}p^{m_{u}}\,,$
where all $m_{i}\leqslant M-1$ and $u\leqslant u^{*}$, then
$p^{m}\beta/r^{*}\leqslant p^{m+M-1}(b_{1}+\dots+b_{u})\leqslant p^{m+M-1}(p-1)<p^{m+M}\leqslant qp^{M-1}$
@@
Suppose that $\iota=p^{m}(qp^{M-1}\alpha-(q-1)\beta)\in\mathfrak{A}^{0}$. Then the sum of the “ $p$-digits” $b_{1}+\dots+b_{u}$ of the appropriate $\beta/r^{*}=b_{1}p^{m_{1}}+b_{2}p^{m_{2}}+\dots+b_{u}p^{m_{u}}$ depends only on $\iota$.
@@
$\mathrm{ch}\,\iota:=b_{1}+\dots+b_{u}$
@@
If $\mathrm{ch}\,\iota\geqslant 1$ then $\iota\neq 0$. (Use that $A[u]\cap B_{1}[u]=\emptyset$.)
In the notation from Sect. 2.2 let $\iota=p^{m}(qp^{M-1}\alpha-(q-1)\beta)\in{\mathfrak{A}}^{0}$. By Lemma 2.2, $p^{m}\alpha$ depends only on $\iota$ and can be presented (non-uniquely) in the form $a_{1}p^{n_{1}}+a_{2}p^{n_{2}}+\dots+a_{u}p^{n_{u}}$ where all coefficients $a_{i}\in[0,(p-1)v_{0}))\cap\mathbb{Z}$, $0\leqslant n_{i}\leqslant m$, $n_{1}=m$ and $u\leqslant u^{*}$.

### v5

@@ 1.4. Explicit description of ideals $\mathcal{L}(w)_{W(k)}$
Set $M(\bar{a},s)=\min\{m\in\mathbb{Z}_{\geqslant 0}\ |\ p^{m}[D_{\bar{a}\bar{n}}]\in\mathcal{L}(r(\bar{a})+s)_{W(k)}\}$.
We always assume that $r=r(\bar{a})<p$ because $D_{\bar{a}\bar{n}}\in\mathcal{L}(r)_{W(k)}$.
If $\bar{a}=\bar{0}$ then, clearly, $M(\bar{a},0)=0$ but $M(\bar{0},s)$ are not defined for $s\geqslant 1$.
Assume that $\bar{a}\neq\bar{0}$.
Let $\bar{u}=\bar{u}(\bar{a})$ be the minimal integer such that for all $1\leqslant i\leqslant r$, there are $p$-digits $\theta_{i},\theta_{i}^{\prime}\in[0,p-1]$ such that
$\theta_{i}/p^{\bar{u}}+\theta_{i}^{\prime}/p^{\bar{u}+1}\leqslant a_{i}/v_{0}<\theta_{i}/p^{\bar{u}}+(\theta_{i}^{\prime}+1)/p^{\bar{u}+1}\,,$
Since $\bar{u}$ is minimal, at least one $\theta_{i}\neq 0$.
Set $\bar{\kappa}_{0}=\kappa_{0}(\bar{a})=\sum_{1\leqslant i\leqslant r}\theta_{i}$ and $\bar{\kappa}_{1}=\kappa_{1}(\bar{a})=\sum_{1\leqslant i\leqslant r}\theta^{\prime}_{i}$.
@@
Proof.
Note that the initial $[D_{\bar{a}\bar{n}}]\in\mathcal{L}(r+s)_{W(k)}$, where $s=0$. The most “economical” way to increase $s$ is to multiply $D_{a_{i}n_{i}}$ with $\theta_{i}\neq 0$ by $p^{\bar{u}}$. We can do this $\bar{\kappa}_{0}$ times to use the inequalities $a_{i}/v_{0}\geqslant\theta_{i}/p^{\bar{u}}$. At this stage we use the basic relations from $\mathcal{R}_{a_{i}}^{I}$. If we have not yet reached $\mathcal{L}(p)_{W(k)}$ we can use $\bar{\kappa}_{1}$ multiplications by $p^{\bar{u}+1}$ to increase $s$ by $\bar{\kappa}_{1}$.
If we still are not in $\mathcal{L}(p)_{W(k)}$ then choose $a_{i_{0}}$ with $\theta_{i_{0}}\neq 0$ and proceed with relations $\mathcal{R}_{a_{i_{0}}}(\theta_{i_{0}}-1,\theta^{\prime}_{i_{0}}+2+t,\bar{u})$. This will require the factor $p^{\bar{u}+2}$ at the first step and the factors $p^{\bar{u}+1}$ at all next steps.
It is easy to verify that the results do not depend on a chosen strategy. ∎
@@
If $\bar{u}=0$ then $[D_{\bar{a}\bar{n}}]\in\mathcal{L}(r+\bar{\kappa}_{0})_{W(k)}$.
@@ 1.5. $\mathbb{Z}_{p}$-submodules $\mathcal{L}_{\mathcal{N}}(w)\subset\mathcal{L}(w)$
For $w\geqslant 1$, let $\mathcal{L}_{\mathcal{N}}(w)$ be the closed submodule in $\mathcal{L}$ such that $\mathcal{L}_{\mathcal{N}}(w)_{W(k)}$ is generated by all $p^{m}[D_{\bar{a}\bar{n}}]$ which satisfy the inequality $|\bar{a}|\geqslant v_{0}U(w-1,m)$.
@@
If $p^{m}[[D_{\bar{a}_{1}\bar{n}_{1}}],[D_{\bar{a}_{2}\bar{n}_{2}}]]\in\mathcal{L}_{\mathcal{N}}(w)_{W(k)}$ then for $i=1,2$, there are $m_{i},w_{i}$ such that:
@@
We can assume that (otherwise, $w_{1}+w_{2}$ is not maximal),
$v_{0}U(w_{i},m_{i})>|\bar{a}_{i}|\geqslant v_{0}U(w_{i}-1,m_{i})\,,i=1,2\,.$
Let $v_{0}U(w_{1}+w_{2},m)>|\bar{a}_{1}|+|\bar{a}_{2}|$. Then $U(w_{1}+w_{2},m)>U(w-1,m)$ (use that $|\bar{a}|\geqslant v_{0}U(w-1,m)$) implies $w_{1}+w_{2}\geqslant w$, q.e.d.
@@
Let $m_{1}^{\prime}\leqslant m_{1}$ and $m_{2}^{\prime}\geqslant m_{2}$. Then $v_{0}U(w_{1},m_{1}^{\prime})\geqslant v_{0}U(w_{1},m_{1})>|\bar{a}_{1}|$ and $v_{0}U(w_{2},m_{2}^{\prime})\leqslant v_{0}U(w_{2},m_{2})$.
@@
Clearly, ¯$\bar{\mathcal{L}}(p)=0$ and $D_{0}$ with $D_{an}$, $a<(p-1)v_{0}$, generate ¯$\bar{\mathcal{L}}_{W(k)}$.
@@ 2.1. Parameters $r^{*}$ and $N^{*}$
Introduce the following subset in $\mathbb{Q}$ :
$A[u^{*}]$ is the set of all
$a_{1}p^{n_{1}}+a_{2}p^{n_{2}}+\dots+a_{u^{*}}p^{n_{u^{*}}}\,,$
where $n_{1},\dots,n_{u^{*}}\leqslant 0$, $n_{1}=0$ and all $a_{i}\in[0,w^{*})\cap\mathbb{Z}$.
The set $A[u^{*}]$ is “left-discrete”, i.e. for any $\gamma>0$, there is $\delta(\gamma)>0$ such that if $\alpha\in A[u^{*}]$ and $\alpha<\gamma$ then $\alpha\leqslant\gamma-\delta(\gamma)$. This property was proved in [9], Sect. 2.1. Therefore, for $1\leqslant s<p$, $0\leqslant m<M$, there is a positive infinum $\delta_{0}(s,m)$ of positive values $v_{0}-\alpha/U(s,m)$, where $\alpha$ runs over the set $A[u^{*}]$.
@@
For $m\in\mathbb{Z}_{\geqslant 0}$, denote by $A[u^{*},m]$ the subset of $A[u^{*}]$ consisted of the elements such that all $n_{u}\geqslant-m$.
@@
Introduce the following subset in $\mathbb{Q}$ :
$B[u^{*}]$ is the set of all numbers
$r^{*}(b_{1}p^{m_{1}}+b_{2}p^{m_{2}}+\dots+b_{u^{*}}p^{m_{u^{*}}})\,,$
where all $m_{i}\leqslant M-1$, $m_{1}\geqslant 0$, $b_{i}\geqslant 0$, $b_{1}\neq 0$, $b_{1}+\dots+b_{u^{*}}<p$. (In particular, $0\notin B[u^{*}]$.) For $m\in\mathbb{Z}_{\geqslant 0}$, $B[u^{*},m]$ is the subset of $B[u^{*}]$ consisted of elements such that all $m_{i}\geqslant-m$.
The above sets $A[u^{*}]$ and an analog of the set $B_{1}[u^{*}]:=p^{-(M-1)}B[u^{*}]$ (where we required $m_{1}=0$ instead of $10\leqslant m_{1}\leqslant M-$ 1) were introduced in [12]. Proceeding similarly to [12] we obtain that $A[u^{*}]\cap B_{1}[u^{*}]=\emptyset$ and there is $\rho(u^{*})>0$ such that
$\rho(A[u^{*}],B_{1}[u^{*}]):=\inf\{|\alpha-\beta|\ |\ \alpha\in A[u^{*}],\beta\in B_{1}[u^{*}]\}\geqslant\rho(u^{*})\,.$
@@
a) Use condition C2). Let $\beta_{1}=\beta/p^{M-1}\in B_{1}[u^{*}]$. Then $q\,|q\alpha-(q-1)\beta_{1}|=q^{2}|\alpha-\beta_{1}+\beta_{1}/q|\geqslant q^{2}|\alpha-\beta_{1}|-\beta_{1}q\geqslant q^{2}\rho(A[u^{*}],\,B_{1}[u^{*}])$ $-r^{*}(p-1)q\geqslant 2r^{*}(p-1)q-r^{*}(p-1)q=r^{*}(p-1)q>b^{*}(p-1)$.
It remains to multiply this inequality by $p^{M-1}$.
b) Use condition C3): $\alpha/U(s,m)<v_{0}$ implies that $\alpha/U(s,m)\leqslant v_{0}-\delta_{0}$ and $q\alpha/U(s,m)\leqslant q(v_{0}-\delta_{0})<qr^{*}(1-q^{-1})=b^{*}$. ∎
@@ 2.2. The set $\mathfrak{A}^{0}$
b) Any $p^{m}\alpha$ appears in the form $a_{1}p^{n_{1}}+\ldots+a_{u^{*}}p^{n_{u^{*}}}$, where all $a_{i}\in[0,(p-1)v_{0})$ and $n_{i}\leqslant m$.
c) All $p^{m}\beta\neq 0$ appear in the form $r^{*}(b_{1}p^{m_{1}}+\ldots+b_{u^{*}}p^{m_{u^{*}}})$, where all $m_{i}\geqslant 0$ and $11\leqslant b_{1}+\ldots+b_{u^{*}}\leqslant p-$ 1.
@@
Suppose $\iota=p^{m}(qp^{M-1}\alpha-(q-1)\beta)\in\mathfrak{A}^{0}$. Then:
@@
c) the integers $p^{m}\alpha$ and $p^{m}\beta/r^{*}$ do not depend on the presentation of $\iota$ in the form $p^{m}(qp^{M-1}\alpha-(q-1)\beta)$ from the above definition of $\mathfrak{A}^{0}$.
@@
a) If $\iota\in\mathfrak{A}^{0}_{0}$ then $\iota=qp^{M-1}(p^{m}\alpha)\in\mathfrak{A}^{0}$ implies that $p^{m}\alpha/(p-1)\leqslant b^{*}/q=r^{*}(1-q^{-1})\in(v_{0}-\delta_{0},v_{0})\,$ and $\mathfrak{A}^{0}_{0}\subset\{qp^{M-1}a\ |\ a\in[0,(p-1)v_{0})\cap\mathbb{Z}\}\,.$ The opposite embedding is obvious.
@@
c) If $\iota=p^{m^{\prime}}(qp^{M-1}\alpha^{\prime}-(q-1)\beta^{\prime})$ is another presentation of $\iota$ then $p^{m}\beta/r^{*}$ and $p^{m^{\prime}}\beta^{\prime}/r^{*}$ are non-negative congruent modulo $qp^{M-1}$ integers and the both are smaller than $qp^{M-1}$. Indeed, we have $\beta/r^{*}=b_{1}p^{m_{1}}+b_{2}p^{m_{2}}+\dots+b_{u^{*}}p^{m_{u^{*}}}\,,$ where all $m_{i}\leqslant M-1$. As a result,
$p^{m}\beta/r^{*}\leqslant p^{m+M-1}(b_{1}+\dots+b_{u^{*}})\leqslant p^{m+M-1}(p-1)<p^{m+M}\leqslant qp^{M-1}$
@@
Suppose that $\iota=p^{m}(qp^{M-1}\alpha-(q-1)\beta)\in\mathfrak{A}^{0}$. Then the sum of “ $p$-digits” $b_{1}+\dots+b_{u^{*}}$ of the appropriate $\beta/r^{*}=b_{1}p^{m_{1}}+\dots+b_{u^{*}}p^{m_{u^{*}}}$ depends only on $\iota$.
@@
$\mathrm{ch}\,\iota:=b_{1}+\dots+b_{u^{*}}$
@@
If $\mathrm{ch}\,\iota\geqslant 1$ then $\iota\neq 0$. (Use Prop. 2.1a).)
Let $\iota=p^{m}(qp^{M-1}\alpha-(q-1)\beta)\in{\mathfrak{A}}^{0}$. By Lemma 2.2, $p^{m}\alpha$ depends only on $\iota$ and can be presented (non-uniquely) in the form $a_{1}p^{n_{1}}+\dots+a_{u}p^{n_{u}}$ with non-zero coefficients $a_{i}\in[1,(p-1)v_{0}))\cap\mathbb{Z}$, $n_{i}\leqslant m$, $n_{1}=m$ and $u\leqslant u^{*}$. (We do not count the zero terms.)